OpenStudy (anonymous):

How can I convert this into standard form? :) y=a(x12)^2+25

hero (hero):

If you meant to convert vertex form to standard form, then y=a(x12)^2+25 is not in vertex form because vertex form is \(y = a(x - h)^2 + k\)

OpenStudy (anonymous):

Yeah it was a typo its supposed to be y=a(x-12)^2+25